I. A family buys a house worth $326,000. They pay $75,000 deposit and take a mortgage for the balance at J12=9% p.a. to be amortized over 30 years with monthly payments.

A. Find the value of the mortgage on their house? (1 mark)
B. Find the value of the monthly payment? (3 marks)
C. Find the loan outstanding after making 20 payments? (4 marks)
D. Find the principal repaid in the 21st payment? (5 marks)

II. Fill out the loan amortization schedule provided in the solution template for the first 5 loan payments. What do you notice about the composition of the payment amount? (6 marks)
1
Expert's answer
2020-06-03T18:25:50-0400

I.326 000-75 000=251 000

A.

"2019.60\times360=727056"

interests=727 056-251 000=476056

B.

"251 000=A\times\frac{1-(1.0075)^{-360}}{0.0075}"

A=2019.60

C.

"727056-2019.60\times20=686664" =

D.

"2019.60*21=42411.60"

II




As part of the amount of payments, the principal amount and interest for using the loan